Trois (joyeuses) contributions l’Informatique aux Mathématiques et à la Physique

Résumé L’Informatique Théorique a dû inventer des modèles pour comprendre le comportement des machines (mots, automates ...). Ce faisant, elle a été contrainte de développer des structures nouvelles qui permettent de résoudre certains problèmes issus des autres sciences.

Three (joyful) contributions of Computer Science to Mathematics and Physics

Abstract Theoretical Computer Science needed to develop models in order to understand the behaviour of computers (words, automata ...). Doing this, one had to invent new structures allowing to solve some problems originated from other sciences.